OpenStudy (anonymous):

what is the equation of the line that is passes through the origin and parallel to the line 5x-3x+2=0?

OpenStudy (anonymous):

passes through origin so constant is zero, next parallel to that line, so find the slope!

OpenStudy (anonymous):

oh.. sorry it's 5x-3y+2=0..

OpenStudy (anonymous):

i really need a solution for this problem.. :)

OpenStudy (anonymous):

@kris100 as mashy said find the slope is m=y2-y1/x2-x1 is the formula?

OpenStudy (anonymous):

@ nudnick.. is that you doms?

OpenStudy (anonymous):

estudier answered you nudnick : "For ax+by = c, slope is -a/b"

OpenStudy (anonymous):

@mashy after solving for the slope what is next?

OpenStudy (anonymous):

@nudnick i just solve for the slope.. what should i do next? have you figured it out?

OpenStudy (anonymous):

@kris i am not doms

OpenStudy (anonymous):

@kris i am not sure about it.....this is what i got y=5x/3 please check. i used the solution for a parallel line...

OpenStudy (anonymous):

Once u have the slope use y = m(x-a) +b for slope m and point (a,b)

OpenStudy (anonymous):

@estudier then what solve for parallel line?

OpenStudy (anonymous):

I just answer u. Parallel lines have same slope. Once u have the slope use y = m(x-a) +b for slope m and point (a,b) Origin is (0,0), right?

OpenStudy (anonymous):

if your line pass through the origin, the equation is y = mx, which means if you have found the slope, you've got the solution

OpenStudy (anonymous):

let the line be y=mx since lines are parallel,slopes are equal m=(-b/a) m=-5/-3 =5/3 the line is hence y=(5/3)x 5x-3y=0